# The galaxy-IGM connection in thesan: observability and information content of the galaxy-Lyman- cross-correlation at z6

### v1

@@ 4.3 The relation between GaLCC and mean free path
The GaLCC carries information on the local ionization field around the sources of reionization. In particular, the amplitude of the peak of transmitted flux is indicative of the strength of the ionizing emissivity of galaxies (see Kakiichi etĀ al., 2018) while its position is linked to the size of ionised regions around galaxies. The latter is, essentially, the mean free path of ionizing photons around the galaxies used to compute the GaLCC ().
@@
The connection between the GaLCC peak position and the size of ionised regions, however, is not straightforward. In fact, the mean free path is sensitive to optical depths , while the GaLCC probes regions with . The latter stems from the fact that the GaLCC is measured using the Ly flux, which is completely absorbed whenever the (local) neutral fraction is . In other words, a photon escaping a galaxy will encounter much sooner a patch of the IGM with than accumulating a total optical depth of , simply because it can easily cross the former without being absorbed, but not the latter (on average). Therefore, we expect the GaLCC peak position to trace just the (size of the) innermost part of the ionised bubbles, where the neutral fraction is significantly above . This is further complicated by the evolution of the background UV density field.
@@
In order to shine light on this issue, in Fig.Ā 17 we compare the mean free path (, solid red line) and peak position evolution (diamonds, see AppendixĀ D for details on how this was computed, as well as for a fit to its redshift evolution). The peak positions has a much shallower evolution of the mean free path, which instead shows a rapid change between and as a result of the percolation of ionised bubbles at the end of cosmic reionization. This is a consequence of the fact that the GaLCC probes regions around the (brightest) sources of ionizing photons, and therefore probes the first regions to be reionised, where the mean free path is expected to be the longest. We compare its evolution to the extrapolation of the fit from Worseck etĀ al. (2014), computed from post-reionization mean free path measurements that are therefore affected only by the evolution of density fluctuations in the IGM. While the GaLCC peak position is approximately smaller than the mean free path for the reason outlined above, their redshift evolution is strikingly similar. This suggests that the GaLCC peak probes regions that are fully ionised already, in line with the physical interpretation of the GaLCC (see the first paragraph of e.g.Ā Sec.Ā 3). We also show with green stars the mean free path around galaxies with () computed at , showing. We find consistently that , as expected, and a redshift evolution intermediate between the extrapolation from Worseck etĀ al. (2014) and the one of . The redshift evolution of is however steeper than the one of the GaLCC peak position, indicating that the evolution of ionised bubbles around the selected sources (traced by the local mean free path) is faster than the evolution of the the highly-ionised region probed by the GaLCC.
@@ 5 Discussion and future perspective
5.1 Which surveys can be used to measure the galaxy-Lyman- cross-correlation?
Leveraging our results presented in Sec.Ā 3, we can outline which current and planned surveys might be able to faithfully measure the GaLCC either using available data or with additional observational efforts. In TableĀ 1 we provide a list of ongoing or planned surveys. For each of them we list the expected number of spectra, their spectral resolution, the S/N and the area covered around each line of sight. It reveals that a number of ongoing surveys can potentially be used to greatly extend the measure of the GaLCC. For instance, the WEAVE-QSO survey appears to be a promising candidate to extend observations of this quantity if complemented with observations of the fields around the QSO (at the moment there are no WEAVE-IFU observations planned for these fields; Welsh, private comm.), or by identifying galaxies through their absorption features in the spectra themselves. Similarly, the āHigh-Redshift Sampleā obtained as part of the 4MOST-QSO survey (Merloni etĀ al., 2019) is poised to obtain a large number of spectra of QSOs at intermediate spectral resolution but without information on the surrounding fields. Should this information be available through complimentary surveys, the potential of the GaLCC could be fully exploited. Finally, the combination of DESI and of its legacy survey can deliver an accurate measurement of this cross-correlation, although the impact of the Legacy Surveys completeness for galaxies at has to be assessed before any definitive claim can be made. Overall, our work shine an optimistic light on the possibility to use ongoing or forthcoming surveys for the measurement of the GaLCC. Many of such QSOs are too faint for traditional Ly studies (which require high S/N to resolve individual features, making their observations prohibitively expensive for faint objects), but the resilience of the GaLCC to noise (see Sec.Ā 3.3) entails that they could potentially be used for this measurement.
@@ 5.1 Which surveys can be used to measure the galaxy-Lyman- cross-correlation?
EIGERā | 6 | |||
ASPIREā³ | 25 | |||
WEAVE-QSOā | 0 | |||
4MOST-QSOā | ā | N/A | 0 | |
DESIā” | ā | N/A |
āKashino etĀ al. (2023); Matthee etĀ al. (2023)
ā³Wang etĀ al. (2023)
āhttps://ingconfluence.ing.iac.es/confluence//display/WEAV; as of now, there is no plan to observe the WEAVE-QSO fields with WEAVE-IFU (Welsh, Pieri, private comm.)
ā High-Redshift sample, Merloni etĀ al. (2019), https://www.eso.org/sci/facilities/develop/instruments/4MOST.html
ā” including the Legacy Surveys, Yang etĀ al. (2023)
@@ 5.2 Can we use galaxies as background sources?
Our finding that even noisy spectra, when used in sufficiently large number and/or with sufficiently high spectral resolution, can faithfully recover the cross-correlation signal (Sec.Ā 3.3) opens up the possibility to use galaxy spectra instead of QSO ones as background sources. Thanks to their larger number density, this would open up a swath of new possibilities, including the study of spatial variations in the GaLCC, extending its study to higher redshift and even directly mapping the impact of ionizing photons from individual bright objects. This would represent an extension of the so-called IGM tomography, i.e.Ā the reconstruction of the large-scale post-reionization density field in the IGM using the Ly absorption in the spectra of background galaxies (typically Lyman break galaxies) as an indicator of the density (e.g.Ā Lee etĀ al., 2018; Horowitz etĀ al., 2022; Newman etĀ al., 2020; Ravoux etĀ al., 2020; Kakiichi etĀ al., 2023). Forthcoming instruments like PFS222https://pfs.ipmu.jp/index.html and MOSAIC333https://elt.eso.org/instrument/MOSAIC/ are poised to significantly extend our ability to perform such studies.
Measuring the GaLCC using background galaxies presents a number of challenges, among which stands out the fact that galaxy spectra are much fainter and more intrinsically variable than QSO ones, both conspiring to make their continuum harder to reconstruct, and therefore the measurement of their transmissivity less precise. Importantly, this has the potential to break one key assumption of our results discussed above, namely that the noise is uncorrelated at the pixel level. In order to progress forward without losing generality, we take a simplified approach. Specifically, we assume errors in the spectra reconstruction result in two types of somewhat idealized modification to the intrinsic normalised flux. While this reduces the direct applicability of our results to real observations, overcoming them would require us to make strong assumption on the type of instrument used to collect the galaxy spectra, as well as on the continuum estimation procedure. For this reason, in this work we elect to remain agnostic to such details and provide a proof of concept analysis of such residual correlated noise, which we hope can guide the design of future observations.
In the first type of reconstruction errors, we assume that the reconstructed normalised flux is the superposition of the true normalized flux and a sinusoidal signal. This could result, for example, from unmodeled variability in the galaxy continuum when it is assumed to be a power-law. Starting from synthetic noise-free spectra with resolution , we modify the transmitted flux (where is the observed flux in a pixel of the spectrum and is the estimated continuum in the same pixel) as:
@@
We show the inferred GaLCC for the two types of reconstruction error described in Fig.Ā 18 and Fig.Ā 19. The former shows the impact of correlated noise in the reconstructed continuum by varying the value of , while the latter reports the impact of errors () in the estimated continuum slope. In both Figures, the left and right panels report the results assuming a pixel-level S/N and S/N (i.e.Ā noise free spectrum), respectively. The figures clearly shows that the investigated reconstruction errors do not significantly affect the inferred GaLCC, as long as they are not extreme ( for the correlated noise and for the slope errors).
This finding is very promising, as it seems to indicate that even large errors in the continuum reconstruction (as expected for galaxies) do not prevent the recovery of the intrinsic GaLCC. However, we expect this conclusion to depend, at least quantitatively, on the specific details of the spectra (e.g.Ā spectral resolution, total length) and on the galaxy sample employed. We leave this for a future work, as it requires to perform survey-specific forecasts that would break the generality we strive for in this paper.
@@ 5.3 Extension to higher Lyman series lines
Another possible development of the GaLCC is the replacement of the Ly line with higher-order Lyman-series lines. Thanks to their lower oscillator strengths, they are less easily saturated, enabling their detection in less ionised regions, and thus higher redshift. However, they are more and more easily contaminated by lower-redshift Ly transmission, limiting the spectral region where they can easily and reliably be identified (for example, this region for Lyman- is approximately a third of the one for Ly). Nevertheless, the use of Lyman- flux is increasingly common in IGM studies at the tail end of reionization (e.g.Ā Eilers etĀ al., 2019; Jin etĀ al., 2023; Ding etĀ al., 2024). This is promising to enable higher-redshift studies of the GaLCC. Moreover, thanks to their larger saturation densities, higher-order Lyman series lines enable us to probe closer to the overdensities where galaxies reside.
As a starting point for a future thorough study of this, we present in Fig.Ā 20 the Galaxy-Lyman series cross-correlation at (solid lines) and (dashed lines, offset by 1 for visual clarity) for the Ly (green line), Ly (yellow line), Ly (red line) and Ly (blue line). The curves share the overall shape (discussed in Sec.Ā 3), but with some differences. Higher-order lines show a smaller and closer-to-galaxies excess of transmitted flux, and the latter is suppressed at smaller scales. This is true at both redshifts shown in the Figure (and we have ensured that it is true for all redshift in the range ), although it is more evident at later times. The shift in peak position and in the region of suppressed flux stem simply from thee fact that the Ly, Ly and Ly transmission require progressively larger neutral fractions to be entirely absorbed, which are achieved in progressively smaller regions around galaxies. The peak amplitude, instead, is reduced because the average transmitted flux for higher-order transitions increases more in the ābackgroundā (i.e.Ā far from the galaxies) than close to the sources. This is due to the increased difficulty of boosting the transmitted flux in already highly-ionised regions, as it is the case close to the sources of ionizing photons) with respect to not-so-ionised ones. Overall, our results indicate that higher-order Lyman transitions can extend to higher redshift our ability to study the galaxy-IGM connection through the GaLCC, as well as to probe regions closer to the galaxy. In fact, it is conceivable to combine the cross-correlation of different Lyman-series line to statistically probe the density field around high-redshift galaxies. We leave a detail study of this to a future work.

### v2

@@ 4.3 The relation between GaLCC and mean free path
The GaLCC carries information on the local ionization field around the sources of reionization. In particular, the amplitude of the peak of transmitted flux is indicative of the strength of the ionizing emissivity of galaxies (see Kakiichi et al., 2018) while its position is linked to the size of ionized regions around galaxies. The latter is, essentially, the mean free path of ionizing photons around the galaxies used to compute the GaLCC ().
@@
The connection between the GaLCC peak position and the size of ionized regions, however, is not straightforward. In fact, the mean free path is sensitive to optical depths to ionizing photons , while the GaLCC probes regions with . The latter stems from the fact that the GaLCC is measured using the Ly flux, which is completely absorbed whenever the (local) neutral fraction is . In other words, a photon escaping a galaxy will encounter much sooner a patch of the IGM with than accumulating a total optical depth of , simply because it can easily cross the former without being absorbed, but not the latter (on average). Therefore, we expect the GaLCC peak position to trace just the (size of the) innermost part of the ionized bubbles, where the neutral fraction is significantly below . This is further complicated by the evolution of the background UV density field.
@@
In order to shine light on this issue, in Fig. 17 we compare the mean free path (, solid red line, computed as described in Garaldi et al. 2022) and peak position evolution (diamonds, see Appendix E for details on how this was computed, as well as for a fit to its redshift evolution). The peak positions has a much shallower evolution of the mean free path, which instead shows a rapid change between and as a result of the percolation of ionized bubbles at the end of cosmic reionization. This is a consequence of the fact that the GaLCC probes regions around the (brightest) sources of ionizing photons, and therefore probes the first regions to be reionized, where the mean free path is expected to be the longest. We compare its evolution to the extrapolation of the fit from Worseck et al. (2014), computed from post-reionization mean free path measurements that are therefore affected only by the evolution of density fluctuations in the IGM. While the GaLCC peak position is approximately smaller than the mean free path for the reason outlined above, their redshift evolution is strikingly similar. This suggests that the GaLCC peak probes regions that are fully ionized already, in line with the physical interpretation of the GaLCC (see the first paragraph of e.g. Sec. 3). We also show with green stars the mean free path around galaxies with () computed at , showing. We find consistently that , as expected, and a redshift evolution intermediate between the extrapolation from Worseck et al. (2014) and the one of . The redshift evolution of is however steeper than the one of the GaLCC peak position, indicating that the evolution of ionized bubbles around the selected sources (traced by the local mean free path) is faster than the evolution of the the highly-ionized region probed by the GaLCC. For comparison, we also report the recent measurements of the mean free path from Gaikwad et al. (2023, green squares) and Zhu et al. (2023, purple circles).
4.4 Local gas properties
Finally, we investigate the local gas properties around the sources and their role in setting the GaLCC. To this end, we compute the volume-averaged radial profile of different simulated quantities around all galaxy with stellar mass ). We chose this stellar mass threshold because it provides a visually-cleaner picture, but we have checked that results are robust against it. We show the resulting profiles in Fig. 18, normalised by their asymptotic () value minus one, for the sake of visual clarity. The quantities shown are: Ly transmissivity (black solid line), HI photoionisation rate (, dot-dashed green line), gas temperature (, dotted red line), HI fraction (blue crosses, to be read against the right-hand-side vertical axis), the HI fraction computed only for gas cells with ( to mimic the Ly visibility, since gas less ionised than this threshold will completely suppress the Ly flux), inverse of the recombination rate (, double-dot-dashed purple line). Note that when the quantity investigated is the Ly transmitted flux (i.e. ), we obtain the GaLCC.
The behaviour of the GaLCC and its relation to the local gas properties is cleanly split at its peak location. On larger scales, the Ly transmission excess follows very closely the excess profile, increasing as the distance form the sources decreases. In this region, the HI recombination rate is approximately constant. Therefore it does not affect the modulation of the Ly transmission with spatial scales. The situation dramatically changes at the GaLCC peak scale. At this and smaller scales, the recombination rate sharply increases (thus its inverse, shown in the plot, decreases). This is a because at these scales the environment of the overdensities sourcing the GaLCC (see Sec. 4.1) becomes important and the enhanced gas density boosts recombination into HI, therefore suppressing the H II fraction and thus the Ly flux.
This competition between the HI photoionization rate and recombination rate effectively modulates the amount of extremely ionised hydrogen, which determines the Ly flux. We stress that, given the large oscillator strength of Ly, only gas with local neutral fraction , since the gas is otherwise fully opaque to Ly regardless of the gas conditions. We explicitly show this in Fig. 18 by plotting two version of the profile around galaxies. One is computed using all gas (blue crosses, to be read against the right-hand-side vertical axis) and shows that the H II fraction is monotonically decreasing with distance, but remains above 90% until , showing no correlation with the GaLCC. However, when all the Ly opaque gas is masked out (by only considering gas with local neutral fraction , shown using a blue dashed line and to be read against the left-hand-side verical axis) the profile is strikingly similar to the GaLCC. Note that we have rescaled this curve to visually match the GaLCC in amplitude. This is necessary because the variations in the H II fraction are tiny (approximately ) compared to those in the Ly flux due to the very large oscillator strength of this line and would not be visible in the plot otherwise. This clarifies further that the GaLCC can not be directly compared to ionised bubbles sizes, since the latter are defined typically using variations in ionised fraction. For instance, in the thesan-1 simulation at the redshift investigated the median bubble size is Neyer et al. (2024), much larger than the GaLCC peak and, in fact, at such scale the Ly transmissivity is already at its average value. It should be noted, however, that box size limitation might quantitatively affect this, since gas quantities are forced to converge to their average value at scales , while the bubble size is computed along individual lines of sight, that do not suffer from this limitation. This has also been hinted at recently in (Conaboy et al., 2025).
We also investigate the role of gas temperature in shaping the GaLCC. The radial profile of this quantity shows two salient features. At scales larger than few cMpc/, it shows a modulation somewhat similar to the GaLCC, although the peak is much smaller in amplitude and at somewhat larger scales. At smaller scales, it exhibits a sharp increase that reaches orders of magnitude above the asymptotic value. This is a consequence of feedback processes within the galaxies that inject large amounts of heat in their surroundings. For the same reason, the H II fraction is also boosted by similar amounts. Overall, these discrepancies points to a negligible impact of gas temperature in setting the GaLCC at this redshift. We verify this by producing synthetic spectra with gas temperature fixed to K and find virtually no difference in the GaLCC. Note, however, that might play a larger role at lower redshift, as reported recently by (Conaboy et al., 2025).
@@ 5 Discussion and future perspective
5.1 Which surveys can be used to measure the galaxy–Lyman- cross-correlation?
Leveraging our results presented in Sec. 3, we can outline which current and planned surveys might be able to faithfully measure the GaLCC either using available data or with additional observational efforts. In Table 1 we provide a list of ongoing or planned surveys. For each of them we list the expected number of spectra, their spectral resolution, the S/N and the area covered around each line of sight. It reveals that a number of ongoing surveys can potentially be used to greatly extend the measure of the GaLCC. For instance, the WEAVE-QSO survey appears to be a promising candidate to extend observations of this quantity if complemented with observations of the fields around the QSO (at the moment there are no WEAVE-IFU observations planned for these fields; Welsh, private comm.), or by identifying galaxies through their absorption features in the spectra themselves. However, its sensitivity may not be sufficient to detect galaxies at the redshift discussed in this paper. Similarly, the ‘High-Redshift Sample’ obtained as part of the 4MOST-QSO survey (Merloni et al., 2019) is poised to obtain a large number of spectra of QSOs at intermediate spectral resolution but without information on the surrounding fields. Should this information be available through complimentary surveys, the potential of the GaLCC could be fully exploited. Finally, the combination of DESI and of its legacy survey can deliver an accurate measurement of this cross-correlation, although the impact of the Legacy Surveys completeness for galaxies at has to be assessed before any definitive claim can be made.
Overall, our work shines an optimistic light on the possibility to use ongoing or forthcoming surveys for the measurement of the GaLCC. Many of such QSOs are too faint for traditional Ly studies (which require high S/N to resolve individual features, making their observations prohibitively expensive for faint objects), but the resilience of the GaLCC to noise (see Sec. 3.3) entails that they could potentially be used for this measurement. In addition, a number of new-generation instruments are currently being deployed or developed. For instance MOONS on VLT (Cirasuolo et al., 2014), SUBARU-PFS (Sugai et al., 2012) and MOSAIC on the future ELT (Hammer et al., 2014) have promising capabilities for observing the GaLCC, thanks to their high multiplexing and resolving power. In the more distant future, an instrument like the Wide-field Spectroscopic Telescope (Mainieri et al., 2024) could make the GaLCC a routine observation.
@@ 5.1 Which surveys can be used to measure the galaxy–Lyman- cross-correlation?
| EIGER∗ | 6 | |||
| ASPIRE△ | 25 | |||
| WEAVE-QSO∘ | 0 | |||
| 4MOST-QSO† | – | N/A | 0 | |
| DESI‡ | – | N/A |
∗Kashino et al. (2023); Matthee et al. (2023)
△Wang et al. (2023)
∘https://ingconfluence.ing.iac.es/confluence//display/WEAV; as of now, there is no plan to observe the WEAVE-QSO fields with WEAVE-IFU (Welsh, Pieri, private comm.)
†High-Redshift sample, Merloni et al. (2019), https://www.eso.org/sci/facilities/develop/instruments/4MOST.html
‡ including the Legacy Surveys, Yang et al. (2023)
@@ 5.2 Can we use galaxies as background sources?
Our finding that even noisy spectra, when used in sufficiently large number and/or with sufficiently high spectral resolution, can faithfully recover the cross-correlation signal (Sec. 3.3) opens up the possibility to use galaxy spectra instead of QSO ones as background sources. Thanks to their larger number density, this would open up a swath of new possibilities, including the study of spatial variations in the GaLCC, extending its study to higher redshift and even directly mapping the impact of ionizing photons from individual bright objects. This would represent an extension of the so-called IGM tomography, i.e. the reconstruction of the large-scale post-reionization density field in the IGM using the Ly absorption in the spectra of background galaxies (typically Lyman break galaxies) as an indicator of the density (e.g. Lee et al., 2018; Horowitz et al., 2022; Newman et al., 2020; Ravoux et al., 2020; Kakiichi et al., 2023). Forthcoming instruments like PFS333https://pfs.ipmu.jp/index.html and MOSAIC444https://elt.eso.org/instrument/MOSAIC/ are poised to significantly extend our ability to perform such studies.
Measuring the GaLCC using background galaxies presents a number of challenges, among which stands out the fact that galaxy spectra are much fainter and more intrinsically variable than QSO ones, both conspiring to make their continuum harder to reconstruct, and therefore the measurement of their transmissivity less precise. Importantly, this has the potential to break one key assumption of our results discussed above, namely that the noise is uncorrelated at the pixel level. In order to progress forward without losing generality, we take a simplified approach. Specifically, we assume errors in the spectra reconstruction result in two types of somewhat idealized modification to the intrinsic normalised flux. While this reduces the direct applicability of our results to real observations, overcoming them would require us to make strong assumption on the type of instrument used to collect the galaxy spectra, as well as on the continuum estimation procedure. For this reason, in this work we choose to remain agnostic to such details and provide a proof of concept analysis of such residual correlated noise, which we hope can guide the design of future observations.
In the first type of reconstruction errors, we assume that the reconstructed normalised flux is the superposition of the true normalized flux and a sinusoidal signal. This could result, for example, from unmodeled variability in the galaxy continuum when it is assumed to be a power-law. Starting from synthetic noise-free spectra with resolution (taken as representative of the spectral resolution of the MOONS, PFS and MOSAIC instruments), we modify the transmitted flux (where is the observed flux in a pixel of the spectrum and is the estimated continuum in the same pixel) as:
@@
We show the inferred GaLCC for the two types of reconstruction error described in Fig. 19 and Fig. 20. The former shows the impact of correlated noise in the reconstructed continuum by varying the value of , while the latter reports the impact of errors () in the estimated continuum slope. In both Figures, the left and right panels report the results assuming a pixel-level S/N and S/N (i.e. noise free spectrum), respectively. The figures clearly show that the investigated reconstruction errors do not significantly affect the inferred GaLCC, as long as they are not extreme ( for the correlated noise and for the slope errors).
This finding is very promising, as it seems to indicate that even large errors in the continuum reconstruction (as expected for galaxies) do not prevent the recovery of the intrinsic GaLCC. However, we expect this conclusion to depend, at least quantitatively, on the specific details of the spectra (e.g. spectral resolution, total length) and on the galaxy sample employed. We leave this for a future work, as it requires to perform survey-specific forecasts that would break the generality we strive for in this paper.
@@ 5.3 Extension to higher Lyman series lines
Another possible development of the GaLCC is the replacement of the Ly line with higher-order Lyman-series lines. Thanks to their lower oscillator strengths, they are less easily saturated, enabling their detection in less ionized regions, and thus higher redshift. However, they are more and more easily contaminated by lower-redshift Ly transmission, limiting the spectral region where they can easily and reliably be identified (for example, this region for Lyman- is approximately a third of the one for Ly). Nevertheless, the use of Lyman- flux is increasingly common in IGM studies at the tail end of reionization (e.g. Eilers et al., 2019; Jin et al., 2023; Ding et al., 2024). This is promising to enable higher-redshift studies of the GaLCC. Moreover, thanks to their larger saturation densities, higher-order Lyman series lines enable us to probe closer to the overdensities where galaxies reside.
As a starting point for a future thorough study of this, we present in Fig. 21 the Galaxy–Lyman series cross-correlation at (solid lines) and (dashed lines, offset by 1 for visual clarity) for the Ly (green line), Ly (yellow line), Ly (red line) and Ly (blue line). The curves share the overall shape (discussed in Sec. 3), but with some differences. Higher-order lines show a smaller and closer-to-galaxies excess of transmitted flux, and the latter is suppressed at smaller scales. This is true at both redshifts shown in the Figure (and we have ensured that it is true for all redshift in the range ), although it is more evident at later times. The shift in peak position and in the region of suppressed flux stem simply from the fact that the Ly, Ly and Ly transmission require progressively larger neutral fractions to be entirely absorbed, which are achieved in progressively smaller regions around galaxies. The peak amplitude, instead, is reduced because the average transmitted flux for higher-order transitions increases more in the ‘background’ (i.e. far from the galaxies) than close to the sources. This is due to the increased difficulty of boosting the transmitted flux in already highly-ionized regions, as it is the case close to the sources of ionizing photons) with respect to not-so-ionized ones. Overall, our results indicate that higher-order Lyman transitions can extend to higher redshift our ability to study the galaxy–IGM connection through the GaLCC, as well as to probe regions closer to the galaxy. In fact, it is conceivable to combine the cross-correlation of different Lyman-series line to statistically probe the density field around high-redshift galaxies. We leave a detail study of this to a future work.